# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 2.3 Curves on toric surfaces
Assuming (3) and imposing the conditions ${\Phi}(p^{\sigma}_{i})=w^{\sigma}_{i}\in{\boldsymbol{w}}_{\partial}$, $1\leq i\leq n(\sigma)$, for every edge $\sigma\in P^{1}_{\Delta}$, as well as imposing extra $n_{{\operatorname{in}}}$ point constraints ${\boldsymbol{w}}_{{\operatorname{in}}}$ in general position in (*$({\mathbb{C}}^{*})^{2}$, we obtain a finite set.
To prove (3), take a generic element $\xi=[{\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})]$ in an irreducible component of ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$. It is a smooth point of that component, and the germ of ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$ at $\xi$ realizes an equisingular deformation. We estimate the dimension of the tangent space to this germ. Recall that, for an isolated singular point $z$ of a planar curve $C$, center of local branches $Q_{1},...,Q_{s}$ (i.e., irreducible components of the germ $(C,z)$), the following holds (see [10, Theorem 1 and formulas (19), (21), (22)]):
  • the tangent space $T(ES_{z})$ to the equisingular stratum $ES_{z}$ in the base space $B_{z}$ of its versal deformation is contained in
    $\left\{b\in B_{z}\ \Big{|}\ \begin{array}[]{c}{\varphi_{b}\in{\mathcal{O}}_{C,z},\ {\operatorname{ord}}_{Q_{i}}\varphi_{b}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {\qquad\qquad\qquad\qquad\qquad+{\operatorname{ord}}Q_{i}-1,\quad i=1,...,s}\end{array}\right\};$
  • the tangent space $T(ES_{z}(L))$, where $L\supset\{z\}$ is a line, to the equisingular stratum $ES_{z}(L)$ in the base space $B_{z}(L)$ of a versal deformation constrained by the fixed intersection multiplicity with the line $L$ at the singular point, is contained in
    $\left\{b\in B_{z}\ \Big{|}\ \begin{array}[]{c}{\varphi_{b}\in{\mathcal{O}}_{C,z},\ {\operatorname{ord}}_{Q_{i}}\varphi_{b}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})}\\ {\qquad\qquad\qquad\qquad\qquad\quad+(Q_{i}\cdot L)-1,\quad i=1,...,s}\end{array}\right\};$
  • the tangent space $T(ES^{fix}_{z}(L))$, where $L\supset\{z\}$ is a line, to the equisingular stratum $ES^{fix}_{z}(L)$ in the base space $B^{fix}_{z}(L)$ of a versal deformation constrained by the fixed position of the singular point $z$ and the fixed intersection multiplicity with the line $L$ at $z$, is contained in
    $\left\{b\in B_{z}\ \Big{|}\ \begin{array}[]{c}{\varphi_{b}\in{\mathcal{O}}_{C,z},\ {\operatorname{ord}}_{Q_{i}}\varphi_{b}\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j})+{\operatorname{ord}}Q_{i}}\\ {\qquad\qquad\qquad\qquad+(Q_{i}\cdot L)-1,\quad i=1,...,s}\end{array}\right\}.$
Let $C={\Phi}(\widehat{C})$. Combining the above description of the ambient space for the considered tangent spaces and taking into account that the conductor ideal $\operatorname{Ann}({\Phi}_{*}{\mathcal{O}}_{\widehat{C}}/{\mathcal{O}}_{C})$ is given by
$\left\{\varphi\in{\mathcal{O}}_{C,z}\ \big{|}\ {\operatorname{ord}}_{Q_{i}}\varphi\geq 2\delta(Q_{i})+\sum_{j\neq i}(Q_{i}\cdot Q_{j}),\ i=1,...,s\right\},$
we derive that
$\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D)\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right),$ (4)
@@
$\deg D=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)$
$+\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap({\mathbb{C}}^{*})^{2}}\sum_{Q\subset(C,z)}({\operatorname{ord}}Q-1).$
Then,
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D)\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right)=C^{2}-\deg D=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2-2g)$
$+CK_{{\operatorname{Tor}}(P_{\Delta})}+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap({\mathbb{C}}^{*})^{2}}\sum_{Q\subset(C,z)}({\operatorname{ord}}Q-1)$
$(2g-2)+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap({\mathbb{C}}^{*})^{2}}\sum_{Q\subset(C,z)}({\operatorname{ord}}Q-1).$ (5)
Here,
$3\leq\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}\leq\#\Delta.$ (6)
Then, by (5) and by the Riemann-Roch formula, the right-hand side of (4) does not exceed
$12g-2+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}-g+$ 1
@@
Furthermore, if either #$\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}<\#\Delta$, or there is a singular local branch $Q$ of $C$ centered in (*$({\mathbb{C}}^{*})^{2}$, then we get ${\mathcal{M}}^{\prime}_{g,n}(\Delta,{\boldsymbol{w}})=\emptyset$. Thus, we are left with the following case: $\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)=n$, the curve $C$ is immersed in (*$({\mathbb{C}}^{*})^{2}$, and #$\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}=\#\Delta$. Arguing on the contrary, we show that $C$ is smooth along the toric divisors. Indeed, suppose that $C$ has a local singular branch $Q_{0}$ centered at some point $w\in{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$. Fixing the position of the point $w$, we obtain a subfamily ${\mathcal{V}}_{1}\subset{\mathcal{M}}^{\prime}_{g,n}(\Delta)$ of dimension $\geq n-1$. On the other hand, the computations as above yield that
$\dim{\mathcal{V}}_{1}\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right),$
@@
$\deg D_{1}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)+{\operatorname{ord}}Q_{0},$
@@
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right)=2g-2+\#\Delta-{\operatorname{ord}}Q_{0}\leq 2g-2+\#\Delta-2,$
and, by the Riemann-Roch formula, we obtain
$\dim{\mathcal{V}}_{1}\leq\#\Delta+g-1-2=n-2,$
which is a contradiction. Thus, $C$ must be totally immersed. Suppose now that some point $w\in C\cap{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$, is a center of local branches $Q_{1},...,Q_{s}$, $s\geq 2$. Again, fixing the position of the point $w$, we obtain a subfamily ${\mathcal{V}}_{2}\subset{\mathcal{M}}^{\prime}_{g,n}(\Delta)$ of dimension $\geq n-1$. On the other hand, the computations as above yield that
$\dim{\mathcal{V}}_{2}\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{2})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right),$
where
$\deg D_{2}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)+s,$
which implies
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-D_{2})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{O}}_{{\operatorname{Tor}}(P_{\Delta})}(C)\right)=2g-2+\#\Delta-s\leq 2g-2+\#\Delta-2,$
@@
It remains to show that ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ does not contain elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ such that ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is a multiple covering of the image $C^{\prime}={\Phi}(\widehat{C})$. Following the preceding argument, we assume that $\dim{\mathcal{M}}_{g,n}(\Delta)=n$ and that a generic element of some component ${\mathcal{V}}_{0}\subset{\mathcal{M}}_{g,n}(\Delta)$ is represented by an $s$-multiple covering ${\Phi}:\widehat{C}\to C^{\prime}\hookrightarrow{\operatorname{Tor}}(P_{\Delta})$, $s\geq 2$, and then come to a contradiction. Denote by ${\Phi}^{\prime}:\widehat{C}^{\prime}\to C^{\prime}$ the normalization and notice that the genus $g^{\prime}$ of $\widehat{C}^{\prime}$ is at most $g$, and that the number of local branches of $C^{\prime}$ centered on the toric divisors is bounded from above by #$\#\Delta$. Further on, we can assume that the variation of ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ along ${\mathcal{V}}_{0}$ induces an equisingular deformation of $C^{\prime}$ in the linear system $|C^{\prime}|$. The computations as in the preceding paragraph imply that
@@
Moreover, the argument of the preceding paragraph would imply that $C^{\prime}$ is immersed and is smooth along the toric divisors. That means that ${\Phi}:\widehat{C}\to\widehat{C}^{\prime}$ is ramified at each of the #$\#\Delta\geq 3$ points of $\widehat{C}^{\prime}$ mapped to the toric divisors, and the ramification index at each of them equals $s$, but this contradicts the Riemann-Hurwitz formula. $\Box$
@@
The set of equivalence classes of such real maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$ considered up to equivariant isomorphism is denoted by ${\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$. We say that an element $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$ is separating if the complement in $\widehat{C}$ to the real points set ${\mathbb{R}}\widehat{C}$ is disconnected, i.e., consists of two connected components. The choice of one of these halves induces the so-called complex orientation on ${\mathbb{R}}\widehat{C}$, as well as on ${\Phi}({\mathbb{R}}\widehat{C})$ if ${\Phi}$ is birational onto its image. Denote by $\overrightarrow{\mathcal{M}}_{g,n}^{{\mathbb{R}}}(\Delta,{\boldsymbol{w}})$ the set of separating elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$ equipped with a choice of a half $\widehat{C}_{+}$ of $\widehat{C}\setminus{\mathbb{R}}\widehat{C}$.
@@
with $z_{1},z_{2}$ coordinates in the torus (×$({\mathbb{C}}^{\times})^{2}\simeq{\operatorname{Tor}}(P_{\Delta})^{\times}$ such that the form $dx_{1}\wedge dx_{2}$ agrees with the orientation of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ defined in Section 2.1. By [19, Theorem 3.1], if all intersection points of ${\Phi}(\widehat{C})$ with the toric divisors are real, then ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$ and $|{\operatorname{QI}}(\xi)|\leq{\mathcal{A}}(\Delta)$.

### v2

@@ 2.4 Curves on toric surfaces
Assuming (4) and imposing the conditions ${\Phi}(p^{\sigma}_{i})=w^{\sigma}_{i}\in{\boldsymbol{w}}_{\partial}$, $1\leq i\leq n(\sigma)$, for every edge $\sigma\in P^{1}_{\Delta}$, as well as imposing extra $n_{{\operatorname{in}}}$ point constraints ${\boldsymbol{w}}_{{\operatorname{in}}}$ in general position in ${\operatorname{Tor}}(P_{\Delta})^{\times}$, we obtain a finite set.
To prove (4), take a generic element $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ in an irreducible component of ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$. It is a smooth point of that component, and the germ of ${\mathcal{M}}^{\prime}_{g,n}(\Delta)$ at $\xi$ realizes an equisingular deformation. We estimate the dimension of the tangent space to this germ. Put $C={\Phi}(\widehat{C})$. Taking into account the formulas in Section 2.3, we derive that
$\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)\leq h^{0}\left(C,{\mathcal{J}}_{Z/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}}\right)=h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right),$ (5)
@@
  • $Z\subset C$ is the zero-dimensional scheme supported at ${\operatorname{Sing}}\hskip 1.70709pt(C)\cup(C\cap{\operatorname{Tor}}(\partial P_{\Delta}))$, defined at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$ by the conductor ideal $I^{cond}(C,z)$, and defined at each point $w\in C\cap{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$, by the ideal described by the right-hand side of the first formula of item (b) in Section 2.3;
  • ${\mathcal{J}}_{Z/C}\subset{\mathcal{O}}_{C}$ is the ideal sheaf of the scheme $Z\subset C$;
  • the divisor ${\boldsymbol{d}}$ satisfies
    $\deg{\boldsymbol{d}}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)$
    $+\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\sum_{Q\subset(C,z)}({\operatorname{mult}}Q-1).$
It follows that
$\deg\left({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right)=C^{2}-\deg{\boldsymbol{d}}$
$=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2-2g)$
$+CK_{{\operatorname{Tor}}(P_{\Delta})}+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}$
$-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\sum_{Q\subset(C,z)}({\operatorname{mult}}Q-1)$
$=(2g-2)+\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}$
$-\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\sum_{Q\subset(C,z)}({\operatorname{mult}}Q-1),$ (6)
where the symbol ##\## stands for the number of elements of a given set. Since
$3\leq\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}\leq\#\Delta,$ (7)
by (6) and by the Riemann-Roch formula, the right-hand side of (5) does not exceed
$12g-2+\#\big{\{}Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\big{\}}-g+$ 1
@@
Furthermore, if either #$\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}<\#\Delta$, or there is a singular local branch $Q$ of $C$ centered in ${\operatorname{Tor}}(P_{\Delta})^{\times}$, then we get ${\mathcal{M}}^{\prime}_{g,n}(\Delta,{\boldsymbol{w}})=\emptyset$. Thus, we are left with the following case: $\dim{\mathcal{M}}^{\prime}_{g,n}(\Delta)=n$, the curve $C$ is immersed in ${\operatorname{Tor}}(P_{\Delta})^{\times}$, and #$\#\{Q\subset(C,w),\ w\in C\cap{\operatorname{Tor}}(\sigma)\}=\#\Delta$. Now we show that $C$ is smooth along the toric divisors. Suppose that $C$ has a local singular branch $Q_{0}$ centered at some point $w\in{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$. Fixing the position of the point $w$, we obtain a subfamily ${\mathcal{V}}_{1}\subset{\mathcal{M}}^{\prime}_{g,n}(\Delta)$ of dimension $\geq n-1$. On the other hand, the computations as above yield that
$\dim{\mathcal{V}}_{1}\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right),$
where by the second formula of item (b) in Section 2.3,
$\deg{\boldsymbol{d}}_{1}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)+{\operatorname{mult}}Q_{0}.$
This implies that
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right)=2g-2+\#\Delta-{\operatorname{mult}}Q_{0}\leq 2g-2+\#\Delta-2,$
and by the Riemann-Roch formula, we obtain
$\dim{\mathcal{V}}_{1}\leq\#\Delta+g-1-2=n-2,$
which is a contradiction. Thus, $C$ must be immersed. Suppose now that some point $w\in C\cap{\operatorname{Tor}}(\sigma)$, $\sigma\in P^{1}_{\Delta}$, is a center of local branches $Q_{1}$, …, $Q_{r}$, $r\geq 2$. Again, fixing the position of the point $w$, we obtain a subfamily ${\mathcal{V}}_{2}\subset{\mathcal{M}}^{\prime}_{g,n}(\Delta)$ of dimension $\geq n-1$. On the other hand, the computations as above yield that
$\dim{\mathcal{V}}_{2}\leq h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{2})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right),$
@@
$\deg{\boldsymbol{d}}_{2}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+\sum_{\sigma\in P^{1}_{\Delta}}\sum_{w\in C\cap{\operatorname{Tor}}(\sigma)}\sum_{Q\subset(C,w)}((Q\cdot{\operatorname{Tor}}(\sigma))_{w}-1)+r,$
@@
$\deg\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{2})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right)=2g-2+\#\Delta-r\leq 2g-2+\#\Delta-2,$
@@
It remains to show that ${\mathcal{M}}_{g,n}(\Delta,{\boldsymbol{w}})$ does not contain elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]$ such that ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is a multiple covering of the image $C^{\prime}={\Phi}(\widehat{C})$. Following the preceding argument, we assume that $\dim{\mathcal{M}}_{g,n}(\Delta)=n$ and that a generic element of some component ${\mathcal{V}}_{0}\subset{\mathcal{M}}_{g,n}(\Delta)$ is represented by a multiple covering ${\Phi}:\widehat{C}\to C^{\prime}\hookrightarrow{\operatorname{Tor}}(P_{\Delta})$, and then come to a contradiction. Denote by ${\Phi}^{\prime}:\widehat{C}^{\prime}\to C^{\prime}$ the normalization. Notice that the genus $g^{\prime}$ of $\widehat{C}^{\prime}$ is at most $g$, and that the number of local branches of $C^{\prime}$ centered on the toric divisors is bounded from above by #$\#\Delta$. Further on, we can assume that the variation of ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ along ${\mathcal{V}}_{0}$ induces an equisingular deformation of $C^{\prime}$ in the linear system $|C^{\prime}|$. The computations as in the preceding paragraph imply that
@@
Moreover, the argument of the preceding paragraph would imply that $C^{\prime}$ is immersed and smooth along the toric divisors. That means that ${\Phi}:\widehat{C}\to\widehat{C}^{\prime}$ is ramified at each of the #$\#\Delta\geq 3$ points of $\widehat{C}^{\prime}$ mapped to the toric divisors, and the ramification index at each of them is maximal, but this contradicts the Riemann-Hurwitz formula. $\Box$
@@
The set of equivalence classes of such real maps ${\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})$ considered up to equivariant isomorphism is denoted by ${\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$. We say that an element $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$ is separating if the complement in $\widehat{C}$ to the real point set ${\mathbb{R}}\widehat{C}$ is disconnected, i.e., consists of two connected components. The choice of one of these halves induces the so-called complex orientation on ${\mathbb{R}}\widehat{C}$, as well as on ${\Phi}({\mathbb{R}}\widehat{C})$ if ${\Phi}$ is birational onto its image. Denote by $\overrightarrow{\mathcal{M}}_{g,n}^{{\mathbb{R}}}(\Delta,{\boldsymbol{w}})$ the set of separating elements $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}^{\mathbb{R}}_{g,n}(\Delta,{\boldsymbol{w}})$ equipped with a choice of a half $\widehat{C}_{+}$ of $\widehat{C}\setminus{\mathbb{R}}\widehat{C}$.
@@
with $z_{1},z_{2}$ coordinates in the torus (×$({\mathbb{C}}^{\times})^{2}\simeq{\operatorname{Tor}}(P_{\Delta})^{\times}$ such that the form $dx_{1}\wedge dx_{2}$ agrees with the orientation of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ defined in Section 2.1. By [24, Theorem 3.1], if all intersection points of ${\Phi}(\widehat{C})$ with the toric divisors are real, then ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$ and $|{\operatorname{QI}}(\xi)|\leq{\mathcal{A}}(\Delta)$.